CONSIDER A CHIMNEY

Boiler uses 20kg of Air per kg of fuel Calculate the minimum height of chimney required to produce a draught of 30mm of water

The mean temperature of chimney gases is 290°c and that of outside is 27°c
1
Expert's answer
2020-06-29T08:27:56-0400

Given H = ?, T2 = 2900 C = 290 +273 = 563 K; T1 = 270 C = 27 +273 = 300 K; m = 20 kg ; h = 30 mm

Solution

h = 353 H ( 1/T1 - (m+1/mT2)

30 = 353 H ( 1/300 - (20+1)/20*563)

H = 57.88 mm minimum height of chimney